### v2

# A compact Optical Liquid Argon Facility at Roma Tre

Abstract
In this paper we present a compact test facility for the measurement of optical properties of liquid argon as scintillation detector. The setup is under preparation at Roma Tre and it has a volume of 40 L liquid argon, which is liquefied from argon gas with a purity of $\geq 99.9999\%$ vol. To readout the scintillation photons from liquid argon with the highest intensity near 127 nm, we use the vacuum ultraviolet silicon photomultipliers from Hamamatsu. By submerging the photon detectors directly inside the liquid argon, we can eliminate the systematics from the wave length shifter and light guides which have been commonly used to detect the scintillation photons of liquid argon.
1 Introduction
Liquid argon (LAr) has been used in neutrino and dark matter experiments as an active medium because of its excellent properties in charge yield and transport, as well as its capacity as a scintillator [1]. To detect LAr scintillation photons of the vacuum ultraviolet (VUV) wavelength range, standard method uses wavelength shifting materials to produce secondary photons in the visible wavelength range to which typical photomultipliers have the sensitivity.
In recent years Hamamatsu has developed the VUV4-series windowless Silicon Photomultipliers (SiPMs), with an appreciable Photon Detection Efficiency (PDE) about 10–20% [2, 3] down to the 127–128 nm peak wavelength of the argon scintillation photon. Without the wavelength shifting material or the light guide, improvement in the systematic uncertainties is expected in the measurement of the optical properties of LAr. The direct detection of the VUV photons will enable new solution and detector design to use LAr as a scintillation detector.
Within the framework of the LEGEND-200 experiment [4], the LLAMA detector has successfully used the Hamamatsu VUV SiPMs inside the 64 m3 LAr passive shielding volume in the LEGEND-200 to monitor the impurity level [5].
At Roma Tre university, we are preparing the Optical Liquid Argon Facility (OLAF), with a cylindrical volume of 40 L LAr to host VUV SiPMs and photon sources inside. The first of the two main goals of the facility is to characterize the optical properties of LAr as a scintillation detector to be used in the LEGEND-1000 experiment [6] as active shielding. Secondly, the facility is expected to function as a test bench for the R&D of the LAr scintillation detector in general. The compact size of the OLAF setup is important for a fast turnaround time to test different configurations and designs of the measurement. In the next section, we will introduce the experimental setup in detail.
2 Experimental setup
The OLAF setup consists of two main parts, to be introduced in the two subsections that follow. The first part is the cryogenic system to liquefy Ar 6.0 gas   [^2] Argon 6.0 is a term for argon gas with a purity level of $\geq 99.9999\%$ vol, the highest available commercially. and to maintain LAr during measurement using liquid nitrogen (LN2) as coolant. The second part is the readout of the vacuum ultraviolet (VUV) SiPMs. The schematic of the setup is indicated in the P&ID diagram in figure 1.
Figure 1: The P&ID diagram of the OLAF setup. Details in text.
2.1 Cryogenics for LAr
The central part of the cryogenic system is a cylindrical vessel made of stainless steel, about 2 m high, as shown in figure 2. The innermost vessel can host LAr up to 40 L, surrounded by a jacketed volume of about 20 L where the LN2 is filled. An external vacuum layer then thermally isolates the inner volumes from the surroundings. The three layers are welded together at the top part of the vessels, as shown in the photo in figure 2a. A 200 L Dewar is used to continuously provide LN2 to OLAF.
Figure 2: The OLAF setup up at the laboratory of Roma Tre: (a) the main vessel with the LAr surrounded by LN2 and vacuum jacketed vessels; (b) photo showing other auxiliary components for OLAF operation.
The thermally insulated pipes were equipped, as shown in figure 2b, to vent the cold gas to the roof outside the laboratory. The top of the LAr volume is fitted with a transparent sight glass through which the interior can be visually inspected. Temperature sensors placed in multiple vertical locations indicate the liquid level of the LAr and LN2 volumes. The feed-through flanges as shown in figure 2a provide separate sockets for the SiPMs and the temperature sensors.
Liquefaction of argon
When the LN2 volume of OLAF is filled, the thermal contact of the jacketed wall provides a cooling power of a few kilowatts to the argon volume. By controlling the flow of the argon gas, we can produce tens of liters of LAr in a few hours  [^3] This includes the time required to cool the vessels from room temperature with LN2. At 1 atm, the boiling temperature of LN2 (77 K) is lower than the freezing temperature of LAr (83 K). To prevent LAr from freezing, we regulate the pressure of the gas phase of the volume of LN2 with a relief valve that vents at 2.5–3 bar ("R2" in figure 1). In stable conditions, the nitrogen volume is sealed by two electric valves indicated by EV1 and EV2 in figure 1, and the boiling temperature of LN2 stays above 86 K.
As the upper rims of the LAr and LN2 vessels have direct thermal contact with the flange at the top of the OLAF setup, a heating power of a few tens of Watts will warm the inner vessels, causing the LAr to evaporate constantly. For measurements longer than a week, we need to replenish LN2 and Ar 6.0 gas to produce more LAr.
2.2 Photon detector layout and readout
The layout of the SiPMs and the light source structure was designed with the primary goal of measuring the attenuation length and the yield of the LAr scintillation photon. The prototype of the mechanical structure for mounting the SiPMs and the photon source is shown in figure 3. The exterior diameter and the height of the cylindrical tower shown in the photo are adapted to the dimension of the OLAF LAr volume inner diameter and the attenuation length of the LAr scintillation photon of the order of some tens of centimeters. The tower has multiple rings each holding one SiPM, placed between 15 cm and 80 cm from the photon sources at the bottom layer of the tower. In figure 3, the photo on the left shows a prototype holder to host the SiPM (delivered by Hamamatsu with a ceramic housing) and to be fixed to a ring. The SiPMs face downward towards the photon sources, and their positions on the rings are staggered to avoid shadowing the SiPMs on the upper levels of the tower.
Figure 3: Middle: the prototype mechanical tower structure which hosts the SiPMs and the photon sources; left: metallic holder to host a single SiPM and to fix it to the tower; top right: a 3D rendering that shows the radioactive source housing and the three SiPMs defining the trigger; bottom right: prototype to test the mounting of LED source to the source holder.
Photon source and calibration
To generate liquid argon scintillation photons of known intensity, we will use a fully encapsulated Am-241 source. The $\alpha$ particles from the primary alpha decay are shielded by the capsule, and only the accompanying 59.5 keV $\gamma$ (35.9% of the time) propagates in LAr with a mean free path of less than 1.5 cm. Using the coincidence of three SiPMs near the point where the $\gamma$ vanishes, one can obtain a well-defined trigger for the position and time of the generated scintillation photons. The 3D rendering in the top-right of figure 3 illustrates the layout for the scintillation photon trigger. We will also place a green LED on the source holder for preliminary function validation and timing calibration of the SiPMs. The figure 3 bottom-right photo shows the LED glued with optical cement to the prototype holder, before the structure was completely submerged in LAr for testing.
Readout electronics
The power supply and the amplification of the SiPM output are realized with a front-end board developed by Roma Tre for the LEGEND-200 experiment [7]. The module is packaged in the form factor of a single NIM unit, and each unit can be connected to 12 SiPMs. Coaxial cables of 5 m long are used to connect the SiPMs to the front-end board. The bias voltage for each connected SiPM can be set separately, while a universal amplification rate of 40 applies to all SiPM signals. The differential outputs of the front-end board are then connected to a CAEN V2740 [8] digitizer module, which has 64 input channels with a sampling rate of 125 MS/sec.
This commercial digitizer has an FPGA on board and can operate with multiple firmwares provided by CAEN. A common feature for all the firmwares is the capability to define a trigger logic during the run-time between the digitized inputs. This feature allows the triple coincidence trigger logic for the scintillation photons to be defined within the digitizer. When the LED is used as the photon source, the digitizer can take an external trigger, which is synchronized with the digital signal modulating the power supply to the LED. In addition to digitizing the waveform of the front-end board’s output, the pulse shape discrimination firmware also provides the integrated charge values of the input waveform with two different user-defined timing gates. Using these two integrated charge values, we can apply the standard technique of pulse shape discrimination for particle type identification for fast online analysis.
Data acquisition
We have implemented the data acquisition of a single V2740 digitizer unit with the open-source package Maximum Integrated Data Acquisition System (MIDAS) [9] developed at PSI and TRIUMF  [^4] Paul Scherrer Institute, Switzerland; TRIUMF, Canada. The program is compiled on an Alma Linux operating system and uses the C++ library provided by CAEN to communicate with the digitizer.
3 Current status and plans
[…]
Figure 4: A typical waveform of the LAr scintillation photon detected by the VUV SiPM. The digitizer operated at the sampling rate of 8 ns per sample; the vertical axis is in ADC unit.
In the next steps, we will implement the readout for multiple SiPMs and the Am-241 source holder structure with the SiPMs for the triple-coincidence trigger, then we will test all SiPMs and photon sources in LAr. The noise level measurement and the optimization of the bias voltage for the SiPM in LAr will follow, as well as the validation of the performance of the trigger. The final production of the mechanical tower structure in oxygen-free copper shall be delivered in 2026, and it will replace the prototype structure. After validating the readout for all the SiPMs and the trigger performance, we will start the physics measurement on the optical properties of the LAr.
Acknowledgments
In preparing the design of the OLAF setup, we thank the LLAMA team at Technical University of Munich: Patrick Krause, Laszlo Papp, Stefan Schönert, Christoph Vogl, and Mario Schwarz. Special thanks to the mechanical workshop team at INFN Sez. Roma Tre: Gianfranco Paruzza, Gabriele Paruzza, and Alessandro Top, who produced the prototype SiPM tower structure and provided invaluable advice on the final design for commercial production. We also thank the HAMMER team at INFN Sez. Roma1 for the 3D printing service.